# CLIPLoss and Norm-Based Data Selection Methods for Multimodal Contrastive Learning

### v1

@@ Abstract
Data selection has emerged as a core issue for large-scale visual-language model pretaining (e.g., CLIP), particularly with noisy web-curated datasets. Three main data selection approaches are: (1) leveraging external non-CLIP models to aid data selection, (2) training new CLIP-style embedding models that are more effective at selecting high-quality data than the original OpenAI CLIP model, and (3) designing better metrics or strategies universally applicable to any CLIP embedding without requiring specific model properties (e.g., CLIPScore is one popular metric). While the first two approaches have been extensively studied, the third remains under-explored. In this paper, we advance the third approach by proposing two new methods. Firstly, instead of classical CLIP scores that only consider the alignment between two modalities from a single sample, we introduce negCLIPLoss, a CLIP loss-inspired method that adds the alignment between one sample and its contrastive pairs as an extra normalization term for better quality measurement. Secondly, when downstream tasks are known, we propose a new norm-based metric, NormSim, to measure the similarity between pretraining data and target data. We test our methods on the data selection benchmark, DataComp [1]. Compared to the best baseline using only OpenAI’s CLIP-L/14, our methods achieve a 5.3% improvement on ImageNet-1k and a 2.8% improvement on 38 downstream evaluation tasks. Moreover, both negCLIPLoss and NormSim are compatible with existing techniques. By combining our methods with the current best methods DFN [2] and HYPE [3], we can boost average performance on downstream tasks by 0.9%, achieving a new state-of-the-art.
@@ 1 Introduction
Curating large-scale visual-language datasets from web-sourced data has become common for pretraining multi-modal models. However, the quality of these web-curated data pairs remains a critical bottleneck. Research has shown that the choice of dataset significantly impacts model performance, irrespective of the models and training techniques employed [4, 5, 6, 7], and this motivates the development of various data selection strategies. This paper focuses on optimizing subset selection from a fixed data pool to train a CLIP model [4] that achieves superior performance on zero-shot downstream tasks.
@@
Classical methods rely solely on OpenAI’s (OAI) pretrained CLIP model (i.e., a teacher model) and focus on better utilizing the embeddings. The most commonly used one is calculating CLIPScore, which measures the cosine similarity between the visual and language embeddings of the CLIP model for the same sample, to eliminate low-quality data with mismatches between text and image. Other works also leverage heuristic distribution alignment techniques to select samples relevant to downstream tasks, such as image-based filtering [1]. These approaches are generally viewed as providing only limited enhancements. However, we argue that the potential of those embeddings has been heavily under-explored. This work seeks a universal method to better employ any given embeddings, not only from OAI CLIP, but also from other CLIP-style models.
@@
On the other hand, recent leading data filtering methods, instead of focusing on improving embedding utilization stategy itself, mainly follow the other two directions, both employing external resources. They either (1) use external non-CLIP models that aid in data selection, (2) or use external high-quality multi-modal data to train a better CLIP-style embedding model than the original OAI CLIP to filter out low-quality data. Specifically, in the first line of works, HYPE [3] leverages embeddings from hyperbolic models instead of the classical Euclidean-based CLIP to measure how each data point has semantically overlaps with other data points and filters out data with low specificity. T-MARS [8] removes images where the text is the only feature correlated with the caption using FAST [9], an off-the-shelf OCR text detection model. Devil [10] applies fasttext [11] to remove non-English texts and use BLIP-2 [12] model for digit recognition to keep useful images with digits. The second direction, represented by Data Filtering Network (DFN) [2], involves training a new CLIP-style teacher model that uses high-quality datasets like HQITP-350M. Although the embeddings extracted from this model perform worse than the OAI CLIP in downstream tasks, it is particularly good at filtering out low-quality data. Notably, some of these methods can be combined and indeed, merging the selected data from DFN and HYPE achieves current state-of-art as shown in HYPE paper [3].
@@
Previous works mainly focus on improving the CLIP embedding quality or utilizing an external model to do filtering but employ the CLIP embedding in a suboptimal way by only using classical methods like CLIPScore. In contrast, in this work, we focus on improving the filtering methods themselves for any given CLIP embedding. We show that there are universal and more effective strategies for utilizing any CLIP teacher model, regardless of its architecture (e.g., B/32 or L/14) or the dataset it was trained on (e.g., OpenAI-WIT-400M or DFN’s high-quality dataset). These strategies should always be orthogonal to the use of any newly trained CLIP-style models like DFN and might also be compatible with methods using external models like FAST and BLIP-2.
Our Contributions. We propose an alternative to CLIPScores that we call negCLIPLoss that more accurately characterizes data quality. We also introduce a new distribution metric we call the p-Norm Similarity Score (NormSim) when knowledge about downstream tasks is available. Two major observations directly inform our proposals:
  • Firstly, we observe that classical methods measure the quality of a multi-modal sample by computing the cosine similarity between its visual and language embeddings, believing that lower similarity indicates that the text does not match its image part well. However, we find that some less informative samples may have a systematic bias, which leads to higher CLIPScores. For example, the language part containing the word "image" can result in higher similarity with any visual part, even when the text does not accurately describe its image content. Our proposed method negCLIPLoss, inspired by the standard CLIPLoss, normalizes the original CLIPScore by the similarity between a sample and its contrastive pairs. For example, the high score caused by the word "image" is typically consistent across its contrastive pairs, so our adjustment reduces this bias. As we have highlighted, such replacement can be universally applied across different embedding models. See Fig. 2 for illustrations.
  • Secondly, if one has access to examples drawn from the same distribution as the target task, it is natural to assume that this extra knowledge could be leveraged to inform the data filtering process. We propose the NormSim metric to measure the vision similarity between a training sample $x$ and the target task dataset $X^{v}_{\text{target}}\in\mathbb{R}^{n\times D}$ defined as $\|f_{v}(X^{v}_{\text{target}})f_{v}(x^{v})\|_{p}$, where $f_{v}:\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ is the vision encoder of teacher model so that $f_{v}(X^{v}_{\text{target}})\in\mathbb{R}^{n\times d}$, $f_{v}(x^{v})\in\mathbb{R}^{d}$, and $f_{v}(X^{v}_{\text{target}})f_{v}(x^{v})\in\mathbb{R}^{n}$, and $\|\cdot\|_{p}$ is the $p$ norm; effective choices are $p=2$ or $\infty$. Notably, unlike previous ImagetNet-based filtering [1], which tries to keep the training set as diverse as downstream tasks by clustering the training set and finding the nearest neighbor group for every target sample, our method does not explicitly consider the diversity but greedily select examples as long as it is close to any target sample (i.e. select high NormSimscore). Notably, negCLIPLoss and NormSim enjoy complementary effect in data selection. See Fig. 3.
[…]
@@
Moreover, the improvements achieved by our methods are not limited to OAI CLIP-based methods but can also be obtained by combining our methods with advanced models that require external resources. By merging the subset selected by negCLIPLoss and NormSim with the subset selected by current state-of-the-art method “HYPE $\cup$ DFN”, we can further improve it by 0.9% on both ImageNet-1k and on average 38 downstream tasks. Besides, we can also achieve a 0.8% improvement on average 38 tasks over "HYPE $\cup$ DFN" using only the data selected by DFN and our strategies. More importantly, we demonstrate that negCLIPLoss, as a replacement for CLIPScore, can be applied to any other embedding models like OAI-L/14, OAI-B/32, and DFN-B/32, universally boosting performance from 0.4% to 3.0% on an average of 38 tasks. This result is not only technically insightful for understanding the information available in embeddings but also practically significant. Compared to existing methods, our approach saves a significant amount of computational time on both reprocessing and new embedding retraining as shown in Table 4.
@@ 2 Problem Setup
Data Filtering on Multimodal Dataset. We are given a training dataset $D_{\text{train}}=\{x^{v},x^{l}\}$, where $(x^{v},x^{l})\in\mathbb{R}^{D}$ is the image-text (vision-language) training pair. For convenience, we will let superscript $vl$ denote either modality so that, for example, $x^{vl}\in{x^{v},x^{l}}$. Our goal is to identify a subset $S\subset D_{\text{train}}$ that maximizes the zero-shot accuracy of the CLIP model on some downstream tasks when $S$ is used to train the CLIP model.
CLIP score and embedding. Recent efforts, such as LAION [5] and DataComp [1], use OpenAI’s CLIP ViT-L/14 model [4] as a teacher model to obtain quality score. Here we denote this vanilla CLIP model as $\bar{f}_{vl}$. For any pair $x^{vl}$, the model outputs a normalized unit-vector $\bar{f}_{vl}(x^{vl})$. If $X^{vl}:=\{x_{1}^{vl},\ldots,x_{m}^{vl}\}$ denotes a dataset containing $m$ samples, then we define $\bar{f}_{vl}(X^{vl})=[\bar{f}_{vl}(x_{1}^{vl}),\ldots,\bar{f}_{vl}(x_{m}^{vl})]^{\top}\in\mathbb{R}^{m\times d}$ as the embedding matrix. The popular filtering metric “CLIPScore” is defined as $\langle\bar{f}_{v}(x^{v}),\bar{f}_{l}(x^{l})\rangle\in[-1,1]$.
Dataset and model. Here we follow the pipeline of Datacomp [1] to standardize the training and evaluation process. This is a testbed for dataset experiments aiming to open-source and further improve the vanilla CLIP model and is widely adopted in previous data selection papers  [13, 14, 8, 2, 15, 7]. We will give more details in Sec. 4.
Figure 1: Illustration of negCLIPLoss. CLIPScore may underestimate (bottom left) or overestimate (bottom right) the quality of image-text pairs. However, this issue can be mitigated by simply adding a normalization term. negCLIPLoss employs the teacher model to calculate the negative CLIP loss on training data and serves as a more accurate metric. Here, “Top X%” denotes that the score represents the top Xhigh values within the entire dataset.
@@ 3 Data Filtering Strategy
In this section, we introduce a better and statistically interpretable quality metric called negCLIPLoss, which directly replaces the common metric CLIPScore. Fig. 1 illustrates how negCLIPLoss works. This new metric only requires negligible extra computational costs and no additional external data collection costs. As the name suggested, this metric is inspired by the standard CLIP loss used in the actual training process of the teacher CLIP model, which is defined as
@@
Here $B^{*}$ is the random batch where $i$-th sample belongs during a particular training step, and $\tau$ is the learnable temperate parameter. Notably, the teacher loss differs from CLIPScore primarily by a normalization term $\mathcal{R}^{*}$ as follows:
@@
In practice, since the training dataset of teacher CLIP models, like OAI-WIT400M [4], and the actual batch divisions $B^{*}$ is inaccessible, we randomly select $K$ batches from the student model’s training data and use the averaged results from $\{B_{k}\}_{i=1}^{K}$ to estimate the normalization term $\mathcal{R}^{*}$ on $B^{*}$, as shown in Eqn. 2:
@@
Here $\{B_{k}\}_{i=1}^{K}$ are some batches randomly selected from the student model’s training data and $x_{i}\in B_{k},\forall k$. We choose $K=10$ in our experiments, but any sample size larger than 5 is sufficiently stable for estimating the original CLIPLoss (Details in Appendix D.1). Besides, in Sec. 4.3.3 we also show that the computational cost introduced by $\mathcal{R}$ remains negligible compared to other baselines. The temperature $\tau$ and batch size $|B^{*}|$ can be directly obtained from the parameters of the pretrained teacher model, and the algorithm process of calculating negCLIPLoss is in Appendix C.2.
@@
Motivation behind negCLIPLoss. Other existing works also use loss-guided data selection, such as LESS [16] in NLP, CoDis [17] in CV, and RHO [18] in general data scheduling scenarios. However, it is still unclear whether selecting based on teacher loss is suitable for multi-modal contrastive learning. Here we give an affirmative answer as shown in Fig. 2, where we can see negCLIPLoss performs better than or on par with CLIPScore consistently.
To illustrate how teacher loss helps our selection, we demonstrate that the normalization term provided by negCLIPLoss is crucial for correcting the overestimation or underestimation inherent in CLIPScore. A high normalization term implies that either the image embedding, text embedding, or both can easily match multiple contrastive pairs beyond their corresponding counterparts. For example, in the bottom right of Fig. 1, the text containing “Image” or “Photo” can be easily matched with any visual content. Similarly, the image of “verloopring” only contains very simple features and can be matched with many words like “white”, “empty” or “circle”, etc. Consequently, despite a high absolute CLIPScore, the relative negCLIPLoss within its batch can be lower. In contrast, the bottom left features highly specific elements in both text and images, such as "Islands Harbor," "American football", and "sheep at green". These elements are specific and less likely to match with contrastive pairs, resulting in a higher relative negCLIPLoss.
@@ 3.2 NormSim: A New Training-Target Similarity Metric
Specifically, we assume access to the target set of downstream tasks and denote them as $X_{\text{target}}=\{x_{\text{target},(1)},\ldots,x_{\text{target},(m)}\}$, where each $x_{\text{target},(i)}\in\mathbb{R}^{d}$ is i.i.d.-sampled from the target downstream distribution $\mathcal{P}_{\text{target}}$ [^1] Although out-of-distribution tasks like “WILDS” have distribution shift between training data and test data, they still provides useful information of the test data., but without overlapping with the test set. Then, for each training sample $x^{vl}$ and the corresponding target set $X_{\text{target}}$, the NormSim is defined as:
$\small\text{{NormSim}}_{p}(X_{\text{target}},x):=\|\bar{f}_{v}(X_{\text{target}}^{v})\bar{f}_{v}(x^{v})\|_{p}=\left(\sum_{x_{t}\in X_{\text{target}}}\left|\langle\bar{f}_{v}(x_{{t}}^{v}),\bar{f}_{v}(x^{v})\rangle\right|^{p}\right)^{1/p}$ (3)
We select the subset $S$ of size $N$ by greedily choosing the samples with the largest NormSim scores. The choice of the norm type $p$ can be based on the data distribution and training process. In this paper, we consider two instantiations of $p$:
When $p=2$, this data selection method is equivalent to selecting a subset that aligns with the principal components of the target set variance. In other words, it can be regarded as
When $p=\infty$, the distance metric can be regarded as an even more optimistic measure, such that a training sample will be selected if it has high similarity to any target sample. Note that this is different from nearest-neighbor-based method used in image-based filtering [1], where they are trying to find the nearest training sample of every target sample. In this case, it can be regarded as:
@@
In Appendix D.3, we also show that our $\text{NormSim}_{\infty}$ can outperform the nearest neighbor selection on the downstream target tasks. Here, we show an example selected via the $\text{NormSim}_{2}$(ImageNet-1k) in Fig. 3, showing that this vision-target-aware method is complementary to the quality-based one.
Choice of Target Data. In the experiment parts, we try two kinds of target data: training data from ImageNet-1k (1.3M) or training data from all 24 accessible downstream tasks (2.1M) [^2] Here we only use the target data for data selection, instead of training on them. The target dataset is significantly smaller than pretraining set like DataComp-medium (128M) or external datasets like HQITP-350M utilized by DFN [2]. . We denote them as $\text{{NormSim}}_{p}$(IN-1k) and $\text{{NormSim}}_{p}$(Target), respectively.
Necessity of using vision-only information. We use only the visual information $x^{v}$ instead of multi-modal information $x^{vl}$ for measuring similarity. This is because common crawled text often has brief captions, making the OAI CLIP language embedding weaker than its visual embedding model [1, 19, 20, 21]. Consequently, the language part cannot characterize the pre-training and downstream task distribution as well as the visual part. This phenomenon is also observed in Gadre et al. [1], where image-based filtering (select data whose image embeddings are similar to that from ImageNet-1k) outperforms text-based filtering (select data whose captions contain words from ImageNet-21k). More ablation studies are provided in Appendix D.4.
Using proxy when downstream $X_{\text{target}}$ is inaccessible. Surprisingly, we show that the 2-norm can also be used when only the pre-training set is available. In this case, we construct a proxy “target” set from the pre-training set itself. Specifically, let $S_{i}$ be the selected subset at step $i$, then we treat the current $S_{i}$ as the proxy “target” set. To construct the next smaller set, we greedily select the next data batch $S_{i+1}$ satisfying $\operatorname*{arg\,max}_{S_{i+1}\subset S_{i}}\sum_{x\in S}\text{{$\text{NormSim}_{2}$}}(S_{i},x),$ until reaching an N size subset. We call this approach $\text{{NormSim}}_{2}$-D (Dynamic) and will specify the algorithm details in Appendix C.3.
Theoretical justification. Unlike many existing methods that force diversity by selecting training samples around each ${\bm{x}}_{\text{target}}$, our strategy greedily maximizes similarity without considering actual coverage. For the $p=2$ case, we demonstrate that maximizing $\text{NormSim}_{2}$ is optimal under a linear model $\bar{f}_{v}$, as shown in Appendix A. Our theorem also provides error guarantees for noisy embeddings and explains when vision-only embeddings outperform combined vision and language embeddings. Recent work by Joshi et al. [22] provides a similar analysis but focuses on high-quality data and cross-variance between images and texts. This approach is less effective than image-only methods for filtering noisy datasets, as discussed above.

### v2

@@ Abstract
Data selection has emerged as a core issue for large-scale visual-language model pretaining (e.g., CLIP), particularly with noisy web-curated datasets. Three main data selection approaches are: (1) leveraging external non-CLIP models to aid data selection, (2) training new CLIP-style embedding models that are more effective at selecting high-quality data than the original OpenAI CLIP model, and (3) designing better metrics or strategies universally applicable to any CLIP embedding without requiring specific model properties (e.g., CLIPScore is one popular metric). While the first two approaches have been extensively studied, the third remains under-explored. In this paper, we advance the third approach by proposing two new methods. Firstly, instead of classical CLIP scores that only consider the alignment between two modalities from a single sample, we introduce negCLIPLoss, a CLIP loss-inspired method that adds the alignment between one sample and its contrastive pairs as an extra normalization term for better quality measurement. Secondly, when downstream tasks are known, we propose a new norm-based metric, NormSim, to measure the similarity between pretraining data and target data. We test our methods on the data selection benchmark, DataComp (gadre2023datacomp). Compared to the best baseline using only OpenAI’s CLIP-L/14, our methods achieve a 5.3% improvement on ImageNet-1k and a 2.8% improvement on 38 downstream evaluation tasks. Moreover, both negCLIPLoss and NormSim are compatible with existing techniques. By combining our methods with the current best methods DFN (fang2023data) and HYPE (kim2024hype), we can boost average performance on downstream tasks by 0.9%, achieving a new state-of-the-art.
Keywords: contrastive learning, visual-language pretraining, data selection, CLIP
@@ 1 Introduction
Curating large-scale visual-language datasets from web-sourced data has become common for pretraining multi-modal models. However, the quality of these web-curated data pairs remains a critical bottleneck. Research has shown that the choice of dataset significantly impacts model performance, irrespective of the models and training techniques employed (radford2021learning; schuhmann2022laion; cherti2023reproducible; bai2023qwen), and this motivates the development of various data selection strategies. This paper focuses on optimizing subset selection from a fixed data pool to train a CLIP model (radford2021learning) that achieves superior performance on zero-shot downstream tasks.
@@
Classical methods rely solely on OpenAI’s pretrained CLIP model (i.e., a teacher model) and focus on better utilizing the embeddings. The most commonly used one is calculating CLIPScore (hessel2021clipscore), which measures the cosine similarity between the visual and language embeddings of the CLIP model for the same sample, to eliminate low-quality data with mismatches between text and image. Other works also leverage heuristic distribution alignment techniques to select samples relevant to downstream tasks, such as image-based filtering (gadre2023datacomp). These approaches are generally viewed as providing only limited enhancements. However, we argue that the potential of those embeddings has been heavily under-explored. This work seeks a universal method to better employ any given embeddings, not only from OpenAI’s (OAI) CLIP, but also from other CLIP-style models.
@@
On the other hand, recent leading data filtering methods, instead of focusing on improving embedding utilization stategy itself, mainly follow the other two directions, both employing external resources. They either (1) use external non-CLIP models that aid in data selection, (2) or use external high-quality multi-modal data to train a better CLIP-style embedding model than the original OAI CLIP to filter out low-quality data. Specifically, in the first line of works, HYPE (kim2024hype) leverages embeddings from hyperbolic models instead of the classical Euclidean-based CLIP to measure how each data point has semantically overlaps with other data points and filters out data with low specificity. T-MARS (maini2023t) removes images where the text is the only feature correlated with the caption using FAST (chen2021fast), an off-the-shelf OCR text detection model. Devil (yu2023devil) applies fasttext (joulin2016bag) to remove non-English texts and use BLIP-2 (li2023blip) model for digit recognition to keep useful images with digits. The second direction, represented by Data Filtering Network (DFN) (fang2023data), involves training a new CLIP-style teacher model that uses high-quality datasets like HQITP-350M. Although the embeddings extracted from this model perform worse than the OAI CLIP in downstream tasks, it is particularly good at filtering out low-quality data. Notably, some of these methods can be combined and indeed, merging the selected data from DFN and HYPE achieves current state-of-art as shown in HYPE paper (kim2024hype).
@@
In this work, we focus on improving the filtering methods themselves for any given CLIP embedding rather than mainly relying on external models or multi-modal datasets. We show that there are universal and more effective strategies for utilizing any CLIP teacher model, regardless of its architecture (e.g., B/32 or L/14) or the dataset it was trained on (e.g., OpenAI-WIT-400M or DFN’s high-quality dataset). These strategies should always be orthogonal to the use of any newly trained CLIP-style models like DFN and might also be compatible with methods using external models like FAST and BLIP-2.
Our Contributions. We introduce two innovative approaches for data quality characterization in multi-modal contexts: negCLIPLoss and Norm Similarity Score (NormSim). negCLIPLoss addresses the limitations of traditional CLIPScores by normalizing scores using the similarity between a sample and its contrastive pairs, effectively reducing biases such as those introduced by generic terms like "image". This method applies universally across various embedding models (see Fig. 2). On the other hand, NormSim leverages available knowledge about downstream tasks to enhance data filtering. It computes the vision similarity between a training sample and a target dataset using a $p$-norm, where significant parameters are typically $p=2$ or $\infty$. Both metrics, negCLIPLoss and NormSim, complement each other in optimizing data selection for specific tasks (see Fig. 3).
@@
To illustrate the effectiveness of our methods, we use a widely used benchmark DataComp (gadre2023datacomp). We show that, by simply replacing CLIPScores with negCLIPLoss and utilizing NormSim, we can exceed the best OAI-CLIP(L/14)-based baseline by 5.3% on ImageNet-1k and 2.8% on average across 38 downstream tasks, which is similar or even better than the performance achieved by many external-resources-based methods. Moreover, by merging the subset selected by negCLIPLoss and NormSim with the subset selected by current state-of-the-art method “HYPE $\cup$ DFN”, we can further improve it by 0.9% on both ImageNet-1k and on average 38 downstream tasks. We can also achieve a 0.8% improvement on average 38 tasks over "HYPE $\cup$ DFN" using only the data selected by DFN and our strategies. Besides, compared to existing methods, our approach saves a significant amount of computational time on data reprocessing as shown in Table 4.
@@ 2 Problem Setup
Data Filtering on Multimodal Dataset. We are given a training dataset $D_{\text{train}}=\{x^{v},x^{l}\}$, where $(x^{v},x^{l})\in\mathbb{R}^{D}$ is the image-text (vision-language) training pair. Our goal is to identify a subset $S\subset D_{\text{train}}$ that maximizes the zero-shot accuracy of the CLIP model on some downstream tasks when $S$ is used to train the CLIP model. CLIP. We denote the teacher CLIP model as $\bar{f}_{vl}$. For any pair $x^{vl}$, the model outputs a normalized unit-vector $\bar{f}_{vl}(x^{vl})$. If $X^{vl}:=\{x_{1}^{vl},\ldots,x_{m}^{vl}\}$ denotes a dataset containing $m$ samples, then we define $\bar{f}_{vl}(X^{vl})=[\bar{f}_{vl}(x_{1}^{vl}),\ldots,\bar{f}_{vl}(x_{m}^{vl})]^{\top}\in\mathbb{R}^{m\times d}$. The popular filtering metric “CLIPScore” is defined as $\langle\bar{f}_{v}(x^{v}),\bar{f}_{l}(x^{l})\rangle\in[-1,1]$. Dataset and model. We follow the pipeline of Datacomp (gadre2023datacomp) to standardize the training and evaluation process as previous papers  (nguyen2023improving; maharana2023d2; maini2023t; fang2023data; mahmoud2023sieve; bai2023qwen).More in Sec. 4.
Figure 1: Illustration of negCLIPLoss. CLIPScore may underestimate (bottom left) or overestimate (bottom right) the quality of image-text pairs. However, this issue can be mitigated by simply adding a normalization term like negCLIPLoss. “Top X%” denotes that the score represents the top Xhigh values within the entire dataset.
@@ 3 Data Filtering Strategy
In this section, we introduce a better and statistically interpretable quality metric called negCLIPLoss, which directly replaces the common metric CLIPScore. Fig. 1 illustrates how negCLIPLoss works. As the name suggested, this metric is inspired by the standard CLIP loss used in the actual training process of the teacher CLIP model:
@@
Here $B^{*}$ is the random batch where $i$-th sample belongs during the final training step, and $\tau$ is the learnable temperate parameter. Notably, the teacher loss differs from CLIPScore primarily by a normalization term $\mathcal{R}^{*}$ as follows:
@@
In practice, since the training dataset of teacher CLIP models, like OAI-WIT400M (radford2021learning), and the actual batch divisions $B^{*}$ is inaccessible, we randomly select $K$ batches from the student model’s training data and use the averaged results from $\{B_{k}\}_{i=1}^{K}$ to estimate the normalization term $\mathcal{R}^{*}$ on $B^{*}$, as shown in Eqn. 2:
@@
Here $\{B_{k}\}_{i=1}^{K}$ are some batches randomly selected from the student model’s training data and $x_{i}\in B_{k},\forall k$. We choose $K=10$ in our experiments, but any sample size larger than 5 is sufficiently stable (Appendix E.1). In Appendix D.2 we also show that the computational cost introduced by $\mathcal{R}$ remains negligible compared to other baselines. Temperature $\tau$ and batch size $|B^{*}|$ can be directly obtained from the pretrained teacher model, and the algorithm process of calculating negCLIPLoss is in Appendix D.3.
Motivation behind negCLIPLoss. Existing works like LESS (xia2024less) in NLP, CoDis (Xia_2023_combatting) in CV, and RHO (pmlr-v162-mindermann22a-prioritized) in data scheduling utilize loss-guided data selection. Despite that, the suitability of this approach for multi-modal contrastive learning remains uncertain. We address this by demonstrating that negCLIPLoss consistently performs as well as or better than CLIPScore (see Fig. 2). The reason is that negCLIPLoss incorporates a normalization term that corrects the biases inherent in CLIPScore (Fig. 1). This term adjusts for cases where text or image embeddings match with multiple contrastive pairs, not just their intended counterparts. For instance, terms like "Image" or "Photo" may result in falsely high CLIPScores as they can match various visual contents. Similarly, simple visual elements can align with numerous general descriptors, lowering their relative negCLIPLoss despite high initial CLIPScores. In contrast, specific words, such as "Islands Harbor" and "American football", have fewer potential matches, resulting in a higher relative negCLIPLoss. This illustrates how negCLIPLoss provides a more accurate measure in multi-modal contexts.
@@ 3.2 NormSim: A New Training-Target Similarity Metric
Specifically, we assume access to the target set of downstream tasks and denote them as $X_{\text{target}}=\{x_{\text{target},(1)},\ldots,x_{\text{target},(m)}\}$, where each $x_{\text{target},(i)}\in\mathbb{R}^{d}$ is i.i.d.-sampled from the target downstream distribution $\mathcal{P}_{\text{target}}$ [^1] Although out-of-distribution tasks like “WILDS” have distribution shift between training data and test data, they still provides useful information of the test data., but without overlapping with the test set. Then, for each training sample $x^{vl}$ and the corresponding target set $X_{\text{target}}$, NormSim is defined as:
$\small\text{{NormSim}}_{p}(X_{\text{target}},x):=\|\bar{f}_{v}(X_{\text{target}}^{v})\bar{f}_{v}(x^{v})\|_{p}$ (3)
We select the subset $S$ of size $N$ by greedily choosing the samples with the largest NormSim scores. We consider $p=\infty$ in the main paper and left $p=2$ in Appendix A.1. When $p=\infty$, a training sample will be selected if it has enough high similarity to any target sample. Note that this is different from the nearest-neighbor-based method used in image-based filtering (gadre2023datacomp), where they are trying to find the nearest training sample of every target sample. In detail, $\text{NormSim}_{\infty}$ can be regarded as:
@@
In Appendix E.4, we also show that our $\text{NormSim}_{\infty}$ can outperform the nearest neighbor selection on the downstream target tasks. Here, we show an example selected via the $\text{NormSim}_{2}$(ImageNet-1k) in Fig. 3 (Refer to Appendix F for more visualization.), showing that this vision-target-aware method is complementary to the quality-based one. Please refer to Appendix A.1 for more discussions.